### v1

# Adapting Vision Foundation Models for Robust Cloud Segmentation in Remote Sensing Images

Abstract
Cloud segmentation is a critical challenge in remote sensing image interpretation, as its accuracy directly impacts the effectiveness of subsequent data processing and analysis. Recently, vision foundation models (VFM) have demonstrated powerful generalization capabilities across various visual tasks. In this paper, we present a parameter-efficient adaptive approach, termed Cloud-Adapter, designed to enhance the accuracy and robustness of cloud segmentation. Our method leverages a VFM pretrained on general domain data, which remains frozen, eliminating the need for additional training. Cloud-Adapter incorporates a lightweight spatial perception module that initially utilizes a convolutional neural network (ConvNet) to extract dense spatial representations. These multi-scale features are then aggregated and serve as contextual inputs to an adapting module, which modulates the frozen transformer layers within the VFM. Experimental results demonstrate that the Cloud-Adapter approach, utilizing only 0.6% of the trainable parameters of the frozen backbone, achieves substantial performance gains. Cloud-Adapter consistently attains state-of-the-art (SOTA) performance across a wide variety of cloud segmentation datasets from multiple satellite sources, sensor series, data processing levels, land cover scenarios, and annotation granularities. We have released the source code and pretrained models at https://github.com/XavierJiezou/Cloud-Adapter to support further research.
Index Terms:
Cloud segmentation, vision foundation models, domain adaptation, fine-tuning, remote sensing image processing.
I Introduction
Figure 1: Comparison between (a) the Previous Method and (b) our Cloud-Adapter approach. Unlike (a), where the entire network (e.g., a typical encoder-decoder architecture) is fully trainable, resulting in a large number of parameters and an increased risk of overfitting, (b) our Cloud-Adapter approach leverages a frozen vision foundation model (VFM) combined with a lightweight, trainable adapter. This design preserves generalization and adaptability, enabling efficient learning for cloud segmentation tasks. The frozen VFM extracts generic visual features, while the parameter-efficient adapter facilitates effective transfer learning across remote sensing scenes.
Cloud segmentation, also known as cloud detection, plays a critical role in the interpretation of remote sensing images, with wide applications in climate monitoring [1, 2], environmental protection, disaster warning, and land use [3]. The variations and characteristics of clouds significantly influence the quality of remote sensing images and subsequent analysis results. Therefore, accurate cloud segmentation is essential for extracting meaningful geospatial information and ensuring reliable data processing. However, due to the diversity of cloud shapes, textures, and sizes, coupled with the complexity of imaging conditions, traditional cloud segmentation methods [4, 5, 6, 7] often fail to achieve stable high-accuracy results in dynamic remote sensing images. Consequently, developing a more robust and adaptable cloud segmentation method has become a key research focus in the field of remote sensing.
With the rapid development of deep learning [8], convolutional neural network (ConvNet [9, 10, 11])-based cloud segmentation methods [12, 13, 14, 15, 16, 17, 18] have made remarkable progress, particularly due to their ability to automatically extract high-level features that enhance the accuracy of cloud shape, edge detection, and texture analysis. For instance, MFCNN [14] combines multi-scale convolutional features to enable fine-grained cloud detection, while DABNet [18] introduces the deformable context feature pyramid module to improve the adaptive modeling capability of multi-scale features for high-accuracy cloud detection. Despite these advancements, many methods still face challenges, including strong dependence on training data, high model complexity, and limited adaptability to diverse remote sensing scenes.
Meanwhile, vision foundation models (VFM) [19, 20, 21, 22, 23], which learn task-agnostic representations from large-scale datasets [24], have demonstrated impressive performance across various applications, such as image classification [25], object detection [26], and segmentation [27, 28, 29]. Transformer architectures like ViT [19] and Swin Transformer [20] have become dominant in the VFM landscape, showing superior performance. However, the potential of VFMs for cloud segmentation remains largely unexplored.
Existing cloud segmentation methods [30, 31] primarily rely on complex neural network architectures, most of which require extensive retraining or fine-tuning on large labeled datasets. Although these methods achieve good performance, they suffer from several limitations. First, they are computationally expensive and require a large amount of labeled data, which is often impractical for real-world applications. Second, traditional models are prone to overfitting, particularly when dealing with small datasets or high-resolution remote sensing images. Third, most existing methods [30, 32, 33, 16, 17, 31, 34] have barely leveraged the power of VFM for cloud segmentation, missing the opportunity to exploit their robust transfer learning capabilities with minimal additional parameters.
To address these challenges, we propose a novel cloud segmentation method based on vision foundation models, named Cloud-Adapter (see Fig. 1). This approach freezes the pretrained VFM backbone and incorporates a lightweight adaptive module, requiring minimal additional parameters to enhance cloud segmentation accuracy and robustness. Specifically, we design a spatial perception module that employs convolutional neural networks (ConvNets) to extract dense spatial features and aggregate them into multi-scale contextual information. This information is then passed into the adaptive module, which modulates the frozen transformer layers of the VFM to efficiently perform cloud segmentation. Experimental results demonstrate that Cloud-Adapter achieves significant performance improvements with only 0.6% of the trainable parameters of the backbone across various popular datasets.
The contributions of this work are summarized as follows:
  1. 1.
    We propose Cloud-Adapter, a simple but powerful parameter-efficient fine-tuning method for cloud segmentation that freezes the backbone of VFMs and fine-tunes only a small fraction of the frozen parameters.
  2. 2.
    We introduce two lightweight components: the spatial perception module and the adapting module, which modulate the features of the VFM, enabling adaptation from a general domain to a remote sensing domain.
  3. 3.
    We demonstrate that our method consistently achieves state-of-the-art performance on four datasets (six subsets) including multiple satellite sources, sensor types, and land cover scenarios, highlighting its robustness towards remote sensing cloud segmentation tasks.
Figure 2: Detailed network architecture of the proposed Cloud-Adapter method, consisting of the spatial perception and adapting modules. The spatial perception module uses ConvNet blocks to extract dense spatial features, which are aggregated into a multi-scale context and fed to the adapting module. The adapting module modulates the frozen transformer layers in the VFM.
II Related Work
II-A Cloud Segmentation
Early cloud segmentation methods primarily relied on the spectral and statistical features of images, using techniques like thresholding, multispectral fusion, and edge detection to distinguish clouds from other surface elements. For instance, CFMask [6, 7] employs decision trees to label pixels and projects estimated cloud heights onto the ground to generate a cloud shadow mask. HOT [4, 5] introduces a transformation optimized for detecting cloud and haze distributions in Landsat images, amplifying haze variations via spectral sensitivity.
However, these traditional methods are highly sensitive to cloud shapes, lighting conditions, and atmospheric environments, resulting in suboptimal performance in complex remote sensing scenarios. With advancements in deep learning, ConvNet-based cloud segmentation methods have gained significant attention due to their ability to automatically extract high-level features, enhancing the accuracy of cloud shape, texture, and edge detection. Examples of innovative model architectures include MSCFF [12], which combines multi-scale convolutional features to facilitate cloud detection. DBNet [31] employs a hybrid model integrating Transformers and ConvNets to capture semantic and spatial details, thereby reducing false detections and omissions in cloud segmentation. HRCloudNet [34] adopts a hierarchical integration approach to preserve cloud texture in high-resolution images, while MCDNet [30] achieves effective thin cloud segmentation through multi-scale feature fusion and auxiliary cloud removal.
For models designed for specific scenarios and training needs, CDNetv1 [16] leverages a feature pyramid and a boundary refinement block to extract cloud masks from low-resolution thumbnails. CDNetv2 [17] improves performance in mixed cloud-snow scenes using adaptive feature fusion and high-level semantic guidance. SCNN [32] applies a shallow ConvNet with filters of varying sizes to minimize model training costs. Additionally, TransMCD [35] uses pseudo-labels derived from spectral features, enabling weakly supervised cloud detection and enhancing model generalization.
Despite these advancements, achieving robust segmentation performance across diverse remote sensing images remains a significant challenge. Furthermore, most methods continue to rely on traditional network architectures, such as UNet [36], ResNet [37], FPN [38], and DeepLab [39], with limited exploration of vision foundation models for cloud segmentation.
II-B Vision Foundation Models
Learning task-agnostic pre-trained representations [40, 41] has become a standard practice in natural language processing. Inspired by this paradigm, substantial progress has been achieved in exploring vision foundation models (VFMs). Typically, these models are trained on large-scale datasets to learn general-purpose representations, achieving remarkable performance across diverse applications, including image classification, object detection, and segmentation. Transformer architectures [19, 20] dominate the landscape of vision foundation models, demonstrating superior performance. CLIP [42] aligns text and image modalities in latent space through contrastive learning. MAE [21] learns robust representations via self-supervised masked modeling. SAM [22] incorporates interactive prompts to capture general concepts of objects, while DINOv2 [23] employs self-distillation to learn generic semantic features. These models effectively capture complex visual patterns and exhibit strong capabilities for few-shot and even zero-shot transfer, allowing them to adapt well to new tasks. Given the diversity and complexity of remote sensing images, exploring vision foundation models with strong generalization is essential for achieving robust cloud segmentation.
II-C Transfer Learning and Domain Adaptation
In resource-constrained environments or when dealing with large-scale data, training models from scratch is often impractical and costly. Transfer learning [43] enables adapting models trained on large datasets to specific tasks. This approach typically involves fine-tuning pre-trained models on smaller, domain-specific datasets, allowing the model to retain general knowledge while learning domain-specific characteristics. Domain adaptation [44] further enhances this process by aligning the feature distributions between the source domain (where the model was initially trained) and the target domain (the new application environment), thus maximizing task performance.
Adapters exemplify the integration of transfer learning and domain adaptation, offering a generally plug-and-play approach across various tasks. Initially introduced in natural language processing (NLP) [45], chapters have rapidly evolved. In NLP tasks, LoRA [46] injects trainable low-rank decomposition matrices into each layer of the Transformer architecture, enabling large language models to adapt to specific domains. In multimodal generation tasks, IP-Adapter [47] is an effective adapter for implementing image prompting functions in pre-trained text-to-image diffusion models. T2I-Adapter [48] introduces a lightweight adapter to modulate the intensity of internal knowledge and external control signals in text-to-image models. For visual segmentation tasks, ViT-Adapter [49] integrates image-related inductive biases into pre-trained ViT models, improving dense prediction performance. SAM-Adapter [50] introduces a simple adapter that incorporates domain-specific information or visual prompts into SAM [22], enhancing performance in specific scenarios. Therefore, developing an efficient adapter shows great potential for improving the robustness of cloud segmentation by leveraging the strengths of vision foundation models.
III Method
In this section, we introduce a simple but powerful fine-tuning method called Cloud-Adapter (as shown in Fig. 2) for cloud segmentation by freezing the vision foundation model (VFM) and incorporating a lightweight adapting module. The method consists of two key components: spatial perception and adapting modules. The spatial perception module extracts multi-scale spatial features using multiple ConvNet blocks and then aggregates them into a dense context. The adapting modules modulate the frozen VFM layers with the context, enabling effective transfer learning for cloud segmentation.
III-A Spatial Perception Module
The spatial perception module (SPM) extracts multi-scale spatial features from the input remote sensing images. First, the input image goes through a stem module for preliminary feature extraction. The stem consists of two depthwise separable convolutions [51]. The first convolution uses a $1\times 1$ kernel to map the input image’s channels to a larger feature space, primarily for channel-wise information fusion. The second convolution uses a $3\times 3$ kernel to extract spatial features and capture local spatial information.
Let the input image be $\mathbf{x}\in\mathbb{R}^{C\times H\times W}$, where $C$ is the number of channels, and $H$ and $W$ are the height and width of the image. The output feature map after stem is $\mathbf{F}_{\text{stem}}\in\mathbb{R}^{C^{\prime}\times H\times W}$, where $C^{\prime}$ is the mapped number of channels, and $H$ and $W$ are the feature map dimensions, which remain the same as the input image.
After the stem, the input features are processed by multiple ConvNet blocks, each of which extracts spatial features at different scales, resulting in a series of feature maps $\mathbf{F}_{i}^{c}\in\mathbb{R}^{C_{i}\times H\times W}$, where $i$ denotes different scales. To aggregate these multi-scale feature maps into a unified scale, we use an aggregator, which applies adaptive average pooling to normalize the feature maps to the same size. The aggregator is a parameter-free operation, meaning it does not contain any trainable parameters, making it a zero-parameter operation.
The aggregation of multi-scale features is expressed as:
$\mathbf{F}_{\text{agg}}=\sum_{i=1}^{k}\mathcal{P}_{avg}(\mathbf{F}_{i}^{c},\text{output\_size}=\mathbf{F}_{k}^{c}.\text{shape}[-2:]),$ (1)
where $\mathbf{F}_{i}^{c}$ are the feature maps at different scales, and $\mathcal{P}_{avg}$ is the adaptive average pooling operation that pools the feature maps to the smallest size of the final feature map $\mathbf{F}_{k}^{c}$.
III-B Adapting Module
Figure 3: Schematic diagram of the proposed adapting module.
The adapting module, as referred to in Fig. 3, utilizes the cross-attention mechanisms [52, 53, 54, 55] to adjust the frozen VFM’s Transformer layers. Specifically, it takes the context $\mathbf{F}_{\text{agg}}$ from the spatial perception module and interacts with the frozen VFM features, thereby modulating the output of the VFM. The cross-attention mechanism calculates the relationships between aggregated dense features and the features extracted by the VFM, producing the adapted features.
The cross-attention computation is expressed by:
$S=\text{Softmax}\left(\frac{\mathbf{QK}^{T}}{\sqrt{d}}\right),$ (2)
$\mathbf{O}=\mathbf{S}\cdot\mathbf{V},$ (3)
$\mathbf{F}_{j}^{att}=\mathcal{M}(\mathbf{O})+\mathbf{F}_{j}^{t},$ (4)
where $\mathbf{Q}\in\mathbb{R}^{d}$, $\mathbf{K}\in\mathbb{R}^{d}$, and $\mathbf{V}\in\mathbb{R}^{d}$ are the query, key, and value matrices, respectively. The Softmax operation computes the attention weights $\mathbf{S}$, which are then multiplied with the value matrix $\mathbf{V}$ to obtain the output $\mathbf{O}$. Finally, after the mapping operation (denoted as $\mathcal{M}$), the adapted features $\mathbf{F}_{j}^{att}$ are obtained with a residual connection, where $j$ denotes the $j$-th feature from the frozen VFM transformer layers.
The mapping operation is performed via a multi-layer perceptron (MLP) with the low-rank [46] design without any activation, aimed to achieve feature projection with minimal parameter cost. The mapping process can be expressed as:
$\mathcal{M}(\mathbf{z})=\mathbf{W}_{2}\cdot(\mathbf{W}_{1}\cdot\mathbf{z}),$ (5)
where $\mathbf{W}_{1}\in\mathbb{R}^{d\times r}$ and $\mathbf{W}_{2}\in\mathbb{R}^{r\times d}$ are the weight matrices of the fully connected layers, $\mathbf{z}$ is the input feature, and $r$ is the rank for dimensionality reduction. The smaller the value of $r$, the fewer trainable parameters the adapting module contains.
IV Experiments
IV-A Datasets
We conducted experiments on four popular cloud segmentation datasets, comprising six subsets, from different satellite sources: CloudSEN12_High [56], L8_Biome [57], GF12MS_WHU [35], and HRC_WHU [12], as summarized in Table I. Specifically, CloudSEN12_High contains two data processing levels, L1C and L2A, while GF12MS_WHU consists of data captured by two satellite sensors, GF1 and GF2. Notably, in typical binary cloud segmentation datasets, thin and thick clouds are grouped into a single category, “cloud,” while cloud shadows are excluded from the annotations.
IV-A1 CloudSEN12_High
[56] CloudSEN12 is a global dataset designed for semantic understanding of clouds and their shadows in Sentinel-2 imagery. It addresses the long-standing challenge of accurately characterizing clouds and cloud shadows in the earth observation community. This dataset comprises 49,400 image patches, including Sentinel-2 Level-1C and Level-2A multispectral data, Sentinel-1 synthetic aperture radar data, auxiliary remote sensing products, and meticulously crafted annotations identifying thick clouds, thin clouds, and cloud shadows. CloudSEN12 stands out due to its diverse annotations, scene variability, broad geographic coverage, and metadata complexity, making it an essential resource for benchmarking and improving cloud detection methodologies. Here, we select its high-quality version with four classes of fine-grained annotations for our experiments.
IV-A2 L8_Biome
[57] The L8_Biome dataset is designed to facilitate comparing and validating cloud detection algorithms for operational Landsat data products. This dataset encompasses various scenario types, including Urban, Barren, Forest, Shrubland, Grass/Cropland, Snow/Ice, Wetlands, and Water. Such a variety of scene types enables models to learn more diverse features, handle complex surface backgrounds, and enhance the accuracy of cloud removal tasks across different terrains. The dataset includes manually generated cloud masks to support the validation of cloud cover assessment algorithms, enabling precise computation of cloud cover percentages for each scene. These characteristics make the dataset an essential resource for cloud segmentation. Here, we cropped each tile into 512 $\times$ 512 patches and removed images that did not contain any targets of interest. The remaining data were then split into training, validation, and test sets in a ratio of 6:2:2.
IV-A3 GF12MS_WHU
[35] The GFMS_WHU dataset consists of 8-meter Gaofen1-MS and 4-meter Gaofen2-MS images. Gaofen1 and Gaofen2 belong to the same satellite series but are equipped with different sensors, offering varying spatial resolutions to enhance the model’s generalization capability. These images were collected from diverse regions of China between June 2014 and December 2020, improving the dataset’s ability to support robust model training. To standardize input dimensions, all images were uniformly cropped and padded to a size of 256 $\times$ 256 patches. The dataset is split into a training set and a test set with a ratio of 7:3, providing a balanced framework for model development and evaluation.
IV-A4 HRC_WHU
[12] HRC_WHU is a high-resolution cloud cover validation dataset created by the SENDIMAGE Lab at Wuhan University. It comprises 150 high-resolution images acquired from Google Earth with RGB channels and resolutions varying from 0.5 to 15 meters. These images have been expertly digitized with two mask annotations: clear sky and clouds. Following the original paper, We uniform cut to 256 $\times$ 256 patch size, 120 for training, and 30 for testing.
TABLE I: Summary of information for four popular cloud segmentation datasets. “-” indicates non-existence or not mentioned.
Dataset Name Satellite Source Ground Resolution Training Set Validation Set Testing Set Patch Size Classes
CloudSEN12_High [56] Sentinel 2 10m 8,490 535 975 512$\times$512 4
L8_Biome [57] Landsat 8 30m 7,931 2,643 2,643 512$\times$512 4
GF12MS_WHU [35] Gaofen 1, 2 8m & 4m 20,697 11,645 - 256$\times$256 2
HRC_WHU [12] Google Earth 0.5m-15m 120 - 30 256$\times$256 2
IV-B Implementation Details
To ensure experimental reproducibility, this section provides a comprehensive overview of the hyperparameter configurations used during training and the evaluation metrics applied to assess the experimental results. All experiments were conducted on a workstation equipped with NVIDIA GeForce RTX 3090 GPUs with 24 GB of memory. The operating system was Ubuntu 22.04, and the neural networks were implemented in PyTorch 2.4.1 with CUDA 11.8. The experiments were performed within the MMSegmentation [58] framework (version 1.2.2), a sound semantic segmentation toolbox based on PyTorch. This framework offers an intuitive interface and a variety of pre-built models, facilitating streamlined experimentation. All segmentation heads across the models used in this study were based on Mask2Former [59], a flexible and efficient architecture tailored for image segmentation tasks. Mask2Former’s unified approach to segmentation and detection tasks contributed to consistent and reliable performance across the different models evaluated in these experiments.
IV-B1 Training Settings
We employed a custom learning rate schedule to optimize model performance. The schedule began with a warm-up phase, utilizing a linear scheduler that increased the learning rate from $1\times 10^{-6}$ over the first 1,000 iterations. This was followed by a PolyLR scheduler with an eta minimum of 0.0 and a power of 0.9, applied from iteration 1,000 to 40,000. The training loop was configured to run for a maximum of 40,000 iterations, with validation conducted at intervals of every 4,000 steps. To ensure reproducibility, we set the random seed to 42 and used the AdamW optimizer [60] with an initial learning rate of $1\times 10^{-4}$ and a weight decay of 0.05. The batch size during training was set to 4, and parameters were initialized using the kaiming normal [61].
IV-B2 Evaluation Metrics
To comprehensively evaluate cloud segmentation performance, we adapt several widely recognized metrics: Intersection over Union (IoU), Accuracy (Acc), and Dice Coefficient (Dice), all reported as percentages (%). Furthermore, we provided the proposed method’s trainable parameters and their proportion relative to the frozen VFM backbone, denoted as “*Params”. The IoU quantifies the overlap between the predicted and the ground truth masks:
$\text{IoU}_{i}=\frac{|\text{Pred}_{i}\cap\text{GT}_{i}|}{|\text{Pred}_{i}\cup\text{GT}_{i}|}=\frac{TP_{i}}{TP_{i}+FP_{i}+FN_{i}},$ (6)
where $TP_{i}$, $FP_{i}$, and $FN_{i}$ represent the true positives, false positives, and false negatives for class $i$, respectively. The Acc represents the pixel-wise classification accuracy:
$Acc_{i}=\frac{TP_{i}+TN_{i}}{TP_{i}+TN_{i}+FP_{i}+FN_{i}},$ (7)
where $TN_{i}$ represents the true negatives for class $i$. The Dice measures the similarity between the predicted and ground truth masks, and it is equivalent to the F1-score metric:
$Dice_{i}=\frac{2|\text{Pred}_{i}\cap\text{GT}_{i}|}{|\text{Pred}_{i}|+|\text{GT}_{i}|}=\frac{2TP_{i}}{2TP_{i}+FP_{i}+FN_{i}}.$ (8)
The metrics mAcc, mIoU, and mDice represent the mean values computed for each class, offering insight into the model’s performance at the class level. In contrast, the overall accuracy (aAcc) evaluates the model’s classification performance across the entire dataset. However, aAcc often overlooks minority classes, such as cloud shadow and thin cloud, potentially overestimating the model’s effectiveness. A significant discrepancy between aAcc and mAcc may highlight class imbalance, where dominant classes skew results. To address this, we calculate the average of these metrics across all courses for cloud segmentation, ensuring a balanced evaluation by considering both per-class accuracy and overall pixel-wise accuracy. These metrics are all positive indicators, with higher values signifying better performance, simplifying interpretation, and facilitating model comparison.
IV-C Ablation Studies and Analysis
This section presents a series of comprehensive ablation studies to evaluate the effectiveness of key components in our proposed model using the CloudSEN12_High_L1C dataset. Specifically, we investigate the impact of (1) different variants of the VFM backbone, (2) the ConvNet architectures employed in the SPM, (3) the dimensions of interactive context, (4) the low-rank MLP in the adapting module, (5) the importance of the aggregator, (6) the interaction frequency within the adapting module, and (7) the presence or absence of each component in the Cloud-Adapter. These studies provide a detailed analysis of how each component contributes to cloud segmentation performance, ultimately highlighting the strengths and advantages of our proposed Cloud-Adapter.
IV-C1 Variants of VFM’s Backbone
TABLE II: Cloud segmentation performance comparison of different VFM’s backbone with their multiple variants.
Backbone Variant mIoU mAcc aAcc mDice *Params (M)
Base 72.53 83.29 89.46 83.22 0.92 (1.1%)
Large 73.36 83.90 89.84 83.85 1.83 (0.6%)
SAM [22] Huge 73.70 84.37 89.95 84.07 2.63 (0.4%)
Small 68.12 79.37 87.50 79.84 0.77 (3.4%)
Base 71.95 83.02 89.22 82.82 0.92 (1.1%)
DINOv2 [23] Large 74.18 84.79 90.19 84.46 1.82 (0.6%)
To evaluate the impact of different vision foundation models and their variants on our method’s performance, we conducted comprehensive experiments using SAM [22] and DINOv2 [23] as backbones. Table II compares three available model variants of them.
We observed that as the model scale increased, both SAM and DINOv2 exhibited consistent performance improvements. Among the SAM variants, the Huge version performed the best, achieving 73.70% mIoU, 84.37% mAcc, 89.95% aAcc, and 84.07% mDice. In contrast, DINOv2 (Large) demonstrated superior performance across all metrics, reaching 74.18% mIoU, 84.79% mAcc, 90.19% aAcc, and 84.46% mDice.
More importantly, DINOv2 (Large) was significantly more computationally efficient, requiring only 11 hours and 8 minutes of training time and 14,558M of GPU memory. In comparison, the SAM (Huge) variant took 19 hours and 9 minutes to train, with GPU memory usage of 24,892M, which was substantially higher than DINOv2 (Large).
These results indicate that DINOv2 (Large) outperforms SAM (Huge) across all metrics while being significantly more efficient in terms of training time and memory usage. Therefore, DINOv2 (Large) emerges as the optimal backbone for our method, striking an excellent balance between performance and computational efficiency for cloud segmentation.
IV-C2 ConvNet Architectures of SPM
TABLE III: Comparison of cloud segmentation performance and model parameters for different ConvNet architectures.
ConvNet Block mIoU mAcc aAcc mDice *Params (M)
Transformer-Like [62] 74.08 84.95 90.09 84.39 1.81 (0.6%)
Pure-ConvNet [55] 74.18 84.79 90.19 84.46 1.82 (0.6%)
We investigate the impact of different ConvNet architectures on our method’s performance, as shown in Table III. Specifically, we compare two types of ConvNet designs: a Transformer-Like architecture inspired by [62] and a Pure-ConvNet structure from [55].
The Pure-ConvNet [55] architecture achieves slightly better overall performance with 74.18% mIoU, 90.19% aAcc, and 84.46% mDice, while the Transformer-Like [62] design shows marginally higher mAcc at 84.95%. Both architectures maintain similar parameter efficiency, with approximately 1.8M.
The comparable performance between these two architectures suggests that both designs are viable options for our method. However, considering the slightly higher performance in most metrics and the simpler structure, we adopt the Pure-ConvNet architecture as our default choice.
IV-C3 Dimension of Interactive Context
We conduct experiments to explore the impact of different dimension settings on the interactive context extracted by the SPM. As illustrated in Fig. 4, we evaluate four configurations: 16, 32, 64, and 128.
The results show that the model’s performance does not necessarily improve with larger dimensions. Starting from dimension 16 with a mIoU of 73.72%, we observe a slight improvement to 73.77% when increasing to dimension 32. The peak performance is achieved at dimension 64 with a mIoU of 74.18%. However, further increasing the dimension to 128 leads to a slight performance degradation (0.44% dropping).
Meanwhile, the number of trainable parameters grows steadily with the dimension size. The model requires 1.60M parameters at dimension 16, gradually increasing to 1.65M at dimension 32, 1.82M at dimension 64, and significantly jumping to 2.49M at dimension 128. Based on these observations, we identify dimension 64 as the optimal choice, offering the best trade-off between performance and efficiency.
Figure 4: Ablation study of different dimension settings on the interactive context extracted by the spatial perception module.
Figure 5: Study of the low-rank MLP in the adapting module.
IV-C4 Low-Rank MLP of Adapting Module
We investigate the influence of different rank settings on our model’s performance and parameter efficiency. The rank dimension controls the size of the low-rank matrices used to approximate the weight updates in the MLP of the adapting module, which helps to reduce the number of trainable parameters. As shown in Figure 5, we examine four rank settings: 4, 8, 16, and 32.
The experimental results reveal an interesting relationship between rank size and model performance. With a rank of 4, the model achieves a mIoU of 73.84% while requiring only 0.61M trainable parameters. Increasing the rank to 8 improves performance to 74.03% with 1.02M parameters. The peak performance is reached at rank 16 with a mIoU of 74.18%, which requires 1.82M parameters. Further increasing the rank to 32 leads to a slight performance degradation. Based on these observations, we choose rank 16 as our optimal setting. This finding aligns with LoRA’s design principle [46] of achieving efficient adaptation through low-rank approximation.
IV-C5 Importance of the Aggregator
TABLE IV: Cloud segmentation performance comparison of context output strategies for the spatial perception module.
Feature Type mIoU mAcc aAcc mDice *Params (M)
Single-Scale 74.05 84.45 90.19 84.34 1.82 (0.6%)
Multi-Scale 74.10 84.70 90.15 84.40 1.82 (0.6%)
Aggregated 74.18 84.79 90.19 84.46 1.82 (0.6%)
We explore three context output strategies for our SPM: Single-Scale, Multi-Scale, and Aggregated approaches. Table IV presents the comparative results of these strategies while maintaining the same parameter count of 1.82M (0.6% of total parameters).
The “Single-Scale” strategy, which directly utilizes the final feature output from the last ConvNet block, achieves baseline performance with 74.05% mIoU, 84.45% mAcc, and 84.34% mDice. The “Multi-Scale” approach, leveraging multiple scale features without fusion, shows slight improvements with 74.10% mIoU, 84.70% mAcc, and 84.40% mDice.
The “Aggregated” strategy, which fuses multiple scale features into a single output with parameter-free operation, demonstrates the best overall performance. It achieves 74.18% mIoU, 84.79% mAcc, and 84.46% mDice, while maintaining the same aAcc (90.19%) as the “Single-Scale” approach. This superior performance can be attributed to the strategy’s ability to effectively combine and utilize information from different feature scales. Based on these results, we adopt the “Aggregated” feature output strategy in our final model design, as it provides the most effective use of dense multi-scale feature information without introducing additional parameters.
IV-C6 Interaction Frequency of Adapting Module
TABLE V: Comparison of performance at different interaction frequencies ($N$) between the Adapter and Transformer layers.
$N$ mIoU mAcc aAcc mDice *Params (M)
4 73.79 84.22 90.10 84.17 0.47 (0.2%)
8 73.76 84.46 89.99 84.15 0.74 (0.2%)
12 73.75 84.46 89.99 84.13 1.02 (0.3%)
24 74.18 84.79 90.19 84.46 1.82 (0.6%)
In this ablation study, we explore the effect of varying the interaction frequency ($N$) between the Adapter and Transformer layers on the model’s performance. The parameter $N$ indicates how frequently the Adapter interacts with the Transformer layers, with larger values corresponding to more frequent interactions. Table V presents the results of this experiment, evaluating different values of $N$: 4, 8, 12, and 24. Specifically, when $N$ is set to 4, the Adapter interacts with the Transformer layers at indices [0, 6, 12, 18]. The “8” and “12” represent interactions at every 3rd and 2nd layer starting from index 0, respectively.
The results indicate that the optimal performance is achieved when $N$ is set to 24, where the model attains 74.18% mIoU, 84.79% mAcc, 90.19% aAcc, and 84.46% mDice. This configuration outperforms all others, particularly in terms of mIoU and mAcc. Notably, the increase in interaction frequency from 4 to 24 leads to improvements in the model’s ability to capture complex spatial dependencies across layers, significantly boosting the performance metrics.
In contrast, when the interaction frequency is lower ($N$ = 4, 8, or 12), performance remains relatively stable but does not reach the peak achieved by the 24-interaction configuration. Furthermore, while the number of parameters increases with higher values of $N$, the increase is minimal (only 0.2-0.6% in parameters), suggesting that the additional interactions do not incur significant computational overhead, thanks to the parameter-efficient LoRA-MLP design. These results suggest that allowing the Adapter to interact with more Transformer layers improves model performance, with the 24-interaction configuration representing a sweet spot for this balance.
IV-C7 Overall Components in the Cloud-Adapter
TABLE VI: Study of each component in the Cloud-Adapter.
Spatial Perception Module Adapting Module mIoU *Params (M)
Stem ConvNet Block Aggregator
$\times$ $\times$ $\times$ $\times$ 73.10 0.00
$\checkmark$ $\times$ $\times$ $\times$ 73.12 3e-4
$\checkmark$ $\checkmark$ $\times$ $\times$ 73.20 0.20
$\checkmark$ $\checkmark$ $\checkmark$ $\times$ 73.37 0.20
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 74.18 1.82
To evaluate the contribution of each component in our model, we conducted an ablation study by systematically removing or modifying specific modules. The results are presented in Table VI, showing the performance in terms of mIoU and parameter count for different configurations. Notably, in configurations without the adapting module, the features output by the spatial perception module are concatenated with the features from the VFM backbone before being passed to the head for prediction.
In the baseline model, without any components from the spatial perception module or adapting module, the features from the VFM backbone were directly passed to the head for segmentation prediction. This configuration achieved a mIoU of 73.10%. When the stem module was added, the mIoU increased slightly to 73.12%, with a negligible parameter increase ($3\times 10^{-4}$). Including the stem module helps in early spatial feature projection and extraction, providing the model with a rudimentary spatial awareness information.
Further improvement was seen when the ConvNet Block and the stem module were added to the model. This resulted in a mIoU of 73.20% and added 0.20M parameters. The ConvNet Block enhances the model’s ability to extract more complex spatial features, contributing to a slight performance boost. When the aggregator module was incorporated, the mIoU rose to 73.37%, with no additional parameters, indicating that the aggregator effectively combines multi-scale features to improve the model’s overall cloud segmentation performance.
Finally, with the inclusion of the adapting module, the full model achieved the highest mIoU of 74.18% and a total of 1.82M parameters. The adapting module enables fine-tuned integration of the spatial and backbone features, significantly enhancing the model’s performance despite increased parameters. This demonstrates the crucial role of the adapting module in optimizing the feature fusion process. In conclusion, the ablation study confirms that each component contributes incrementally to the model’s performance. The concatenation of spatial and backbone features, followed by successive refinement and adaptation through the added modules, substantially improved the segmentation performance in special domains.
IV-D Comparison with State-of-the-Art Methods
We compare the performance of our proposed Cloud-Adapter with several state-of-the-art cloud segmentation methods across different datasets: HRC_WHU, GF12MS_WHU_GF1, GF12MS_WHU_GF2, CloudSEN12_High_L1C, and CloudSEN12_High_L2A. These datasets include both binary (HRC_WHU, GF1, GF2) and multi-class (L1C, L2A) cloud segmentation. The performance comparison, as shown in Tables VII, VIII, IX, X, and XI, highlights Cloud-Adapter’s robustness in both binary and multi-class cloud segmentation.
IV-D1 Binary-Class Cloud Segmentation
TABLE VII: Cloud segmentation performance comparison of different methods on the HRC_WHU dataset.
Method aAcc mIoU mAcc mDice
MCDNet [30] 75.14 53.50 68.91 67.96
SCNN [32] 74.51 57.22 81.27 72.31
KappaMask [33] 84.73 67.48 80.30 79.74
CDNetv2 [17] 89.71 76.75 87.46 86.46
DBNet [31] 90.11 77.78 88.80 87.17
CDNetv1 [16] 89.88 77.79 89.93 87.20
RSAM-Seg [63] 92.07 80.90 88.72 89.16
UNetMobv2 [56] 92.13 79.91 85.61 88.45
HRCloudNet [34] 92.93 83.44 92.39 90.79
Cloud-Adapter (Ours) 94.50 89.05 93.74 94.19
TABLE VIII: Cloud segmentation performance comparison of different methods on the GF12MS_WHU_GF1 dataset.
Method aAcc mIoU mAcc mDice
SCNN [32] 97.18 81.68 87.21 89.13
CDNetv1 [16] 96.81 81.82 92.75 89.27
CDNetv2 [17] 97.55 84.93 92.96 91.36
MCDNet [30] 97.55 85.16 93.97 91.51
RSAM-Seg [63] 98.36 89.26 94.70 94.09
UNetMobv2 [56] 98.82 91.71 93.99 95.53
DBNet [31] 98.73 91.36 95.19 95.33
HRCloudNet [34] 98.80 91.86 95.82 95.62
KappaMask [33] 98.91 92.42 95.05 95.94
Cloud-Adapter (Ours) 98.92 92.55 95.99 96.02
TABLE IX: Cloud segmentation performance comparison of different methods on the GF12MS_WHU_GF2 dataset.
Method aAcc mIoU mAcc mDice
KappaMask [33] 90.30 72.00 77.64 82.57
HRCloudNet [34] 91.46 75.57 80.95 85.29
SCNN [32] 91.99 76.99 82.06 86.32
CDNetv1 [16] 92.42 78.20 83.07 87.17
MCDNet [30] 92.30 78.36 83.95 87.31
DBNet [31] 92.61 78.68 83.39 87.50
CDNetv2 [17] 92.63 78.84 83.67 87.61
UNetMobv2 [56] 93.22 80.44 84.86 88.70
Cloud-Adapter (Ours) 94.04 83.02 87.54 90.40
Clear Sky Thick Cloud
Figure 6: Comparison of visualized segmentation results of different models on the coarse-grained cloud segmentation dataset.
Quantitative Comparison
For binary-class cloud segmentation datasets, Cloud-Adapter consistently demonstrates superior performance across all evaluation metrics (aAcc, mIoU, mAcc, and mDice). On the HRC_WHU dataset (Table VII), Cloud-Adapter achieves the highest scores with an aAcc of 94.50%, mIoU of 89.05%, mAcc of 93.74%, and mDice of 94.19%, significantly outperforming the second-best method, HRCloudNet, which only achieves a mIoU of 83.44%. On the GF12MS_WHU_GF1 dataset (Table VIII), Cloud-Adapter attains an aAcc of 98.92%, mIoU of 92.55%, mAcc of 95.99%, and mDice of 96.02%, surpassing KappaMask, which achieved a slightly lower mIoU of 92.42%. On the GF12MS_WHU_GF2 dataset (Table IX), Cloud-Adapter maintains its leading performance with an aAcc of 94.04%, mIoU of 83.02%, mAcc of 87.54%, and mDice of 90.40%, showing its robustness in handling challenging satellite data.
Visualization Comparison
Fig. 6 compares the visualized segmentation results for three binary-class datasets. Cloud-Adapter stands out for its ability to accurately capture cloud boundaries and reduce over-segmentation issues, which are common in other methods like DBNet and HRCloudNet. The results clearly illustrate its superiority in segmenting clear sky and cloud regions, even under complex conditions.
IV-D2 Multi-Class Cloud Segmentation
TABLE X: Cloud segmentation performance comparison of different methods on the CloudSEN12_High_L1C dataset.
Method IoU mIoU Acc mAcc Dice mDice aAcc
CRS TNC TKC CDS CRS TNC TKC CDS CRS TNC TKC CDS
SCNN [32] 56.39 00.00 34.60 00.00 22.75 89.69 00.00 45.08 00.00 33.69 72.11 00.00 51.41 00.00 30.88 60.19
KappaMask [33] 76.84 14.75 73.51 00.00 41.27 97.54 27.78 77.73 00.00 50.76 86.91 25.70 84.73 00.00 49.33 76.27
MCDNet [30] 68.39 13.02 65.26 32.52 44.80 77.85 18.39 84.49 59.20 59.99 81.23 23.04 78.98 49.08 58.08 72.68
CDNetv1 [16] 80.48 36.50 79.60 44.83 60.35 90.60 50.19 88.34 60.72 72.46 89.19 53.48 88.64 61.91 73.30 83.48
DBNet [31] 85.20 43.02 81.40 52.44 65.52 95.56 53.15 89.40 62.47 75.15 92.01 60.16 89.75 68.80 77.68 86.83
CDNetv2 [17] 84.98 43.37 80.67 53.40 65.60 94.80 53.13 89.32 65.34 75.65 91.88 60.50 89.30 69.62 77.83 86.68
HRCloudNet [34] 86.01 45.88 83.54 57.61 68.26 94.49 59.47 91.39 67.19 78.13 92.48 62.90 91.03 73.10 79.88 87.86
UNetMobv2 [56] 88.51 50.12 84.79 63.17 71.65 95.77 61.87 91.44 75.13 81.05 93.90 66.78 91.77 77.43 82.47 89.52
Cloud-Adapter (Ours) 89.19 56.15 85.46 65.93 74.18 94.08 75.23 91.72 78.15 84.79 94.29 71.91 92.16 79.47 84.46 90.19
TABLE XI: Cloud segmentation performance comparison of different methods on the CloudSEN12_High_L2A dataset.
Method IoU mIoU Acc mAcc Dice mDice aAcc
CRS TNC TKC CDS CRS TNC TKC CDS CRS TNC TKC CDS
SCNN [32] 62.20 00.00 52.74 00.08 28.76 85.73 00.00 76.19 00.08 40.50 76.70 00.00 69.06 00.16 36.48 67.14
KappaMask [33] 79.58 23.35 78.20 00.00 45.28 97.24 38.05 86.58 00.00 55.47 88.63 37.85 87.77 00.00 53.56 79.60
MCDNet [30] 72.46 14.96 67.04 31.62 40.18 85.76 21.09 85.29 42.01 55.94 84.03 26.02 80.27 48.04 53.07 69.21
CDNetv1 [16] 82.58 38.61 78.72 49.64 62.39 92.73 52.65 88.12 60.50 73.50 90.46 55.71 88.09 66.35 75.15 84.74
DBNet [31] 85.42 42.76 80.80 53.62 65.65 94.83 53.53 90.27 63.45 75.52 92.14 59.90 89.38 69.81 77.81 86.82
CDNetv2 [17] 85.35 44.01 80.85 54.00 66.05 95.65 56.37 88.06 63.56 75.91 92.09 61.12 89.41 70.13 78.19 86.88
HRCloudNet [34] 87.24 44.18 82.78 59.20 68.35 95.41 51.00 93.17 69.83 77.35 93.19 61.28 90.58 74.37 79.85 88.35
UNetMobv2 [56] 87.76 47.38 84.16 62.13 70.36 95.98 59.60 90.94 71.75 79.57 93.48 64.29 91.40 76.64 81.45 88.96
Cloud-Adapter (Ours) 88.99 55.04 85.31 65.64 73.75 94.02 74.93 91.48 77.41 84.46 94.17 71.00 92.07 79.26 84.13 89.99
TABLE XII: Cloud segmentation performance comparison of different methods on the L8_Biome dataset.
Method IoU mIoU Acc mAcc Dice mDice aAcc
CRS TNC TKC CDS CRS TNC TKC CDS CRS TNC TKC CDS
MCDNet [30] 65.83 08.40 60.97 00.21 33.85 73.80 10.26 95.01 00.22 44.82 79.40 15.49 75.75 00.41 42.76 69.75
SCNN [32] 66.30 00.00 63.24 00.00 32.38 81.11 00.00 95.12 00.00 44.06 79.73 00.00 77.48 00.00 39.30 71.22
CDNetv1 [16] 60.69 19.63 56.79 01.21 34.58 88.19 25.51 67.43 01.22 45.59 75.54 32.81 72.44 02.40 45.80 68.16
KappaMask [33] 77.72 22.63 68.12 00.00 42.12 93.08 32.87 80.96 00.00 51.73 87.45 36.91 81.04 00.00 51.35 76.63
UNetMobv2 [56] 79.18 28.35 80.46 03.05 47.76 93.62 34.63 93.59 03.31 56.29 88.37 44.18 89.16 05.92 56.91 82.00
CDNetv2 [17] 73.81 24.85 75.85 00.00 43.63 90.18 31.95 89.81 00.00 52.98 84.93 39.81 86.27 00.00 52.75 78.56
HRCloudNet [34] 72.41 30.89 70.76 00.00 43.51 85.50 43.70 85.88 00.00 53.77 83.99 47.20 82.88 00.00 53.52 77.04
DBNet [31] 80.19 38.18 82.37 04.90 51.41 96.51 49.34 87.89 05.03 59.70 89.00 55.25 90.34 09.34 60.99 83.62
Cloud-Adapter (Ours) 85.81 44.41 70.75 29.13 57.53 90.63 78.76 75.35 32.65 69.34 92.36 61.51 82.87 45.12 70.46 81.79
TABLE XIII: Comparison of cloud segmentation performance across different land cover types in the L8_Biome dataset.
Scene Method mIoU mAcc mDice aAcc
Grass/Crops KappaMask [33] 36.29 62.25 42.17 76.56
CDNetv1 [16] 39.99 59.01 44.22 73.04
CDNetv2 [17] 55.17 71.84 59.23 86.78
HRCloudNet [34] 50.47 66.21 56.19 81.99
MCDNet [30] 28.92 69.85 32.92 81.61
SCNN [32] 43.81 70.48 46.86 83.13
DBNet [31] 59.08 73.85 63.31 89.85
UNetMobv2 [56] 49.46 74.78 54.39 88.21
Cloud-Adapter (Ours) 41.60 63.10 48.12 72.71
Urban KappaMask [33] 48.83 73.39 55.55 86.78
CDNetv1 [16] 46.30 64.64 51.03 81.44
CDNetv2 [17] 56.27 74.55 61.74 89.13
HRCloudNet [34] 56.47 71.57 62.85 87.99
MCDNet [30] 31.90 70.68 36.48 83.87
SCNN [32] 48.63 71.55 52.05 85.66
DBNet [31] 59.68 76.73 65.25 92.14
UNetMobv2 [56] 50.24 79.90 55.39 91.59
Cloud-Adapter (Ours) 51.76 76.84 58.36 86.27
Wetlands KappaMask [33] 40.38 61.72 47.59 75.12
CDNetv1 [16] 34.50 56.45 39.69 66.57
CDNetv2 [17] 40.54 59.62 46.71 74.78
HRCloudNet [34] 41.48 57.89 47.93 76.38
MCDNet [30] 26.23 55.86 30.75 68.51
SCNN [32] 34.38 55.90 37.76 67.42
DBNet [31] 49.79 67.54 56.52 83.53
UNetMobv2 [56] 41.54 64.58 47.82 77.59
Cloud-Adapter (Ours) 50.16 74.26 57.09 86.08
Snow/Ice KappaMask [33] 24.41 43.24 29.37 56.74
CDNetv1 [16] 20.93 38.00 23.82 44.90
CDNetv2 [17] 24.99 41.18 28.51 55.37
HRCloudNet [34] 22.20 36.91 25.41 47.58
MCDNet [30] 10.43 25.64 13.30 29.59
SCNN [32] 12.43 23.33 14.37 24.70
DBNet [31] 29.72 47.25 34.04 67.69
UNetMobv2 [56] 23.50 41.99 27.96 56.48
Cloud-Adapter (Ours) 32.01 56.99 36.25 73.15
Scene Method mIoU mAcc mDice aAcc
Barren KappaMask [33] 38.36 60.64 45.31 76.96
CDNetv1 [16] 31.50 48.05 36.15 59.96
CDNetv2 [17] 41.24 59.24 46.03 76.33
HRCloudNet [34] 44.33 60.53 50.19 77.51
MCDNet [30] 26.56 60.52 31.34 73.00
SCNN [32] 36.93 60.42 40.55 73.35
DBNet [31] 47.41 65.13 53.05 80.38
UNetMobv2 [56] 43.66 67.80 49.67 84.01
Cloud-Adapter (Ours) 48.95 72.35 56.45 83.49
Forest KappaMask [33] 38.32 57.00 44.46 80.07
CDNetv1 [16] 33.86 48.50 38.67 71.81
CDNetv2 [17] 41.94 58.62 47.87 81.34
HRCloudNet [34] 42.93 58.30 49.64 81.76
MCDNet [30] 30.05 57.39 35.23 77.71
SCNN [32] 37.62 55.44 41.39 77.92
DBNet [31] 44.93 62.40 50.81 85.22
UNetMobv2 [56] 45.23 64.89 51.17 86.89
Cloud-Adapter (Ours) 48.09 68.76 56.25 82.30
Shrubland KappaMask [33] 40.87 59.71 48.28 77.26
CDNetv1 [16] 32.70 49.79 38.89 61.35
CDNetv2 [17] 42.04 59.48 48.51 76.94
HRCloudNet [34] 41.68 57.22 48.98 76.79
MCDNet [30] 28.66 57.34 34.20 71.57
SCNN [32] 37.11 56.93 41.61 71.25
DBNet [31] 46.49 63.41 53.76 80.88
UNetMobv2 [56] 41.58 63.24 48.19 79.79
Cloud-Adapter (Ours) 51.18 72.01 59.19 87.42
Water KappaMask [33] 42.42 64.40 49.36 81.19
CDNetv1 [16] 44.43 61.84 49.20 82.83
CDNetv2 [17] 48.72 64.17 54.10 85.07
HRCloudNet [34] 46.74 60.46 52.96 82.62
MCDNet [30] 29.43 60.95 33.81 79.71
SCNN [32] 41.82 61.56 45.51 80.86
DBNet [31] 52.42 67.85 58.29 87.15
UNetMobv2 [56] 49.58 70.25 55.37 88.24
Cloud-Adapter (Ours) 46.43 67.75 53.26 82.05
Clear Sky Thick Cloud Thin Cloud Cloud Shadow
Figure 7: Comparison of visualized segmentation results of different models on the CloudSEN12_High_L1C dataset.
Clear Sky Thick Cloud Thin Cloud Cloud Shadow
Figure 8: Comparison of visualized segmentation results of different models on the CloudSEN12_High_L2A dataset.
Clear Sky Thick Cloud Thin Cloud Cloud Shadow
Figure 9: Comparison of visualized segmentation results of different models on the L8_Biome dataset.
Quantitative Comparison
Cloud-Adapter achieves state-of-the-art performance on multi-class cloud segmentation datasets, which include four categories: Clear Sky (CRS), Thin Clouds (TNC), Thick Clouds (TKC), and Cloud Shadows (CDS). On the CloudSEN12_High_L1C dataset (Table X), Cloud-Adapter achieves a mIoU of 74.18%, with IoUs of 89.19% (CRS), 85.46% (TKC), 56.15% (TNC), and 65.93% (CDS). It outperforms methods like UNetMobv2 and HRCloudNet, which show lower scores, particularly in challenging classes such as TNC and CDS. On the CloudSEN12_High_L2A dataset (Table XI), Cloud-Adapter achieves a mIoU of 73.75%, with IoUs of 88.99% (CRS), 85.31% (TKC), 55.04% (TNC), and 65.64% (CDS). Its mDice score of 84.13% further highlights its balanced performance across all categories. On the L8_Biome dataset (Tables XII and XIII), Cloud-Adapter attains the highest mIoU (57.53%), mAcc (69.34%), mDice (70.46%), and aAcc (81.79%). It excels in challenging categories like Snow/Ice (32.01%) and Shrubland (51.18%), consistently outperforming competitors such as DBNet and UNetMobv2 across diverse land scenes.
Visualization Comparison
Fig. 7 and 8 present Cloud-Adapter’s segmentation results on multi-class datasets (CloudSEN12_High_L1C and CloudSEN12_High_L2A), accurately distinguishing thick clouds, thin clouds, and cloud shadows in complex conditions. The figures also demonstrate its performance across varying cloud cover levels: “Low” (less than 35%), “Medium” (35–65%), and “High” (over 65%). These results highlight Cloud-Adapter’s robustness and adaptability to diverse atmospheric scenarios. Fig. 9 highlights Cloud-Adapter’s segmentation performance on the L8_Biome dataset, demonstrating superior generalization across diverse biomes, such as barren, forest, and snow/ice regions. These results further emphasize the model’s capability to adapt to varying environmental conditions while maintaining high accuracy and precision, solidifying its effectiveness for complex cloud segmentation tasks in remote sensing applications.
In conclusion, across both binary-class and multi-class datasets, Cloud-Adapter consistently outperforms state-of-the-art methods. Its high aAcc, mIoU, mAcc, and mDice scores demonstrate its capability to effectively handle diverse cloud segmentation tasks. The visual comparisons further validate its robustness and adaptability, making it a reliable algorithm for cloud detection in remote sensing applications.
V Conclusion
In this paper, we introduced Cloud-Adapter, a parameter-efficient fine-tuning method for cloud segmentation in remote sensing images. By leveraging a frozen VFM and adding a lightweight adapter, Cloud-Adapter reduces the number of trainable parameters while achieving SOTA performance across various datasets. Our approach offers an effective solution for robust cloud segmentation without requiring extensive retraining, making it simultaneously efficient and adaptable.
Limitations. The performance of Cloud-Adapter depends heavily on the VFM, and the large parameters of VFMs make it difficult to deploy on edge devices with limited resources.